OpenStudy (anonymous):

Complete the square to solve the equation. x^2+x=11/4

OpenStudy (anonymous):

\[x^2+x=\frac{11}{4}\] \[(x+\frac{1}{2})^2=\frac{11}{4}+\frac{1}{4}=\frac{12}{4}=3\] \[(x+\frac{1}{2})^2=3\] \[x+\frac{1}{2}=\pm 3\] \[x=-\frac{1}{2}\pm3\]

OpenStudy (anonymous):

Thanks satellite you're soooo smart (: